# DBCopilot: Scaling Natural Language Querying to Massive Databases

### v2

@@ 3.2. Schema Graph Construction
Unlike NL expressions, a valid SQL query must follow the structure of the database. Therefore, it is necessary to inform the schema router about the inherent relationships/constraints between schema elements. In general, the schema $S=\langle D,T\rangle$ involved in a valid SQL query must satisfy the following conditions. First, the set of tables $T$ must be the subset of tables belonging to database $D$, since a SQL query obviously cannot reference tables that do not exist in the target database. Second, the tables in $T$ should cross-reference at least one other table with a common column (e.g., foreign key), so that the SQL query can link them together via join or subquery clauses. In addition to explicit Primary-Foreign relations, two tables can also be linked by implicit Foreign-Foreign relations, i.e., they reference the same unique key column of another table.
@@
Given a question “Which rivers run through the state with the largest city in the US?”, the corresponding SQL query “SELECT river_name FROM river WHERE traverse IN ( SELECT state_name FROM city WHERE population = ( SELECT MAX ( population ) FROM city ) )” uses only the city and river tables, both of which refer to the primary key of the state table.
To this end, we construct a directed graph $\mathcal{G}=\langle\mathcal{V},\mathcal{E}\rangle$ to represent the underlying schema structure of all databases $\mathcal{D}$ and their tables $\mathcal{T}$. Concretely, the schema graph consists of three types of nodes $\mathcal{V}=\{\nu_{s}\}\cup\mathcal{D}\cup\mathcal{T}$, where $\nu_{s}$ is a special node denoting the set of all databases $\mathcal{D}$. Each edge $e\in\mathcal{E}$ in this graph represents a relation between two nodes, which can be divided into:
  • Inclusion relations: indicate whether a database is part of a database collection and whether a table belongs to a database.
  • Table relations: include explicit Primary-Foreign and implicit Foreign-Foreign relations between table nodes.
@@
Input: A dictionary of all database schemas $databases$.
@@ 4       $links\leftarrow$ a dictionary with empty set as default value
8             for $column$ in $table$ :
9                   if $column$ contains a foreign key :
10                         $ltbl$, $lcol$leftarrow $column.{\mkern-2.0muf\mkern-2.0mu}oreign\_key$
11                         Add $ltbl$ and $table$ into set $links[ltbl.lcol]$
@@
16
17
18      for $tables$ in $links$ :
19             for $tbl1$, $tbl2$ in $tables\times tables$ :
20                   if $tbl1$ is not equal to $tbl2$ :
21                         $\mathcal{G}$.addEdge($tbl1$, $tbl2$)
22
23
24
25
26
27
28
29
30return $\mathcal{G}$
@@
Algorithm 1 shows the procedure for constructing this three-tiered hierarchical schema graph. By capturing the structure of database schemas in graph $\mathcal{G}$, any valid SQL query schema $S$ is a trail (Wikipedia contributors, 2023b) on this graph. Each schema trail starts with node $\nu_{s}$, followed by a database node and some of its interconnected table nodes. The schema graph can efficiently narrow down the schema space $\mathcal{S}$ into trials on this graph starting from node $\nu_{s}$. As a result, the schema graph is essential for effectively representing and sampling SQL query schemas, serving as the foundation of DBCopilot.
@@ 3.3. Schema Serialization
Since we model schema routing as a Seq2Seq generative retrieval process, we need to design a schema serialization algorithm that converts a SQL query schema into a sequence of tokens.
Basic Serialization.  A trivial solution is to randomly order the set of tables as ad hoc retrieval (Chen et al., 2023), and identify the database based on the votes of the generated tables. However, this method loses sight of the inclusion and table relations between schema elements. The former results in a huge search space due to the lack of database constraints during table generation. The latter can result in missing tables during schema routing, especially junction tables.
@@
where $\pi$ is a node iteration order and $\operatorname{DFS_{\pi}}(S,\mathcal{G})$ is the DFS traversal sequence of schema $S$ on schema graph $\mathcal{G}$ when iterating successor nodes under order $\pi$.
We introduce the iteration order parameter $\pi$ to determine the DFS sequence because a SQL query schema is not a completely ordered set and there may be multiple DFS sequences for a schema. We randomly select one DFS sequence at a time as the model target during training and assign different probabilities to different sequences originating from the same schema during inference.
@@
Input: An ordered pair of a SQL query schema $S=\langle D,T\rangle$, the constructed schema graph $\mathcal{G}$, and a node iteration order $\pi$.
@@ 10      $successors$leftarrow [][\;][ ]
             /* Iterates over successor nodes according to node iteration order $\pi$/
@@ 3.4. Training Data Synthesis
Based on the schema graph and the DFS serialization, we can learn the schema router on question-schema pairs ($N$, $S$) as described in § 3.5. However, the generative retrieval paradigm prevents the schema router from generalizing to unseen schemas because they are not injected into the model parameters. To address this challenge, we propose a labor-free training data synthesis method that automatically generates pseudo-instances via a reverse schema-to-question generation paradigm, as shown in Figure 2.
Specifically, we sample a sufficient number of valid schemas from the schema graph and generate a pseudo-question for each sampled schema. These schemas are sampled by performing finite-length random walks over the schema graph, and the database and tables traversed by a walk form a sampled schema. To generate an NL question for each sampled schema, we train a schema questioning model $\mathcal{M}_{\text{q}}$ in reverse based on the training sets of NL2SQL datasets. We use a SQL parser to extract the metadata (tables and columns) of SQL queries from the training pairs ($N$, $Q$). Unlike the schema router, which takes a condensed SQL query schema as output, the schema questioner accepts more detailed schema information (such as table columns) as input to generate rich and detailed questions. Given an illustration of the model input and output in Figure 3, the training loss of the schema questioning model $\mathcal{M}_{q}$ with parameters $\theta_{\text{q}}$ on $m$ extracted schema-question pairs ($S$, $N$) is:
@@
Pseudo-Question Generation Input: Ask a question that needs to be answered by combining the contents of all the tables, based on the word1 database schema provided below. - countrylanguage(language, countrycode, official) - country(code, lifeexpectancy) Output: What is the official language of the country with the highest life expectancy?
@@
3.5. Schema Router Training & Inference
Through the above measures, we can efficiently train the schema router to parameterize the semantic mapping of massive database schemas and effectively route NL questions to their target schemas via graph-based constrained decoding. Formally, the Seq2Seq schema router $\mathcal{M}_{r}$ takes a question sequence $N=\left[t_{1},t_{2},\ldots,t_{|N|}\right]$ as input and generates the serialized schema sequence $\bar{S}=\left[s_{1},s_{2},\ldots,s_{|\mathbf{s}|}\right]$. Suppose the synthetic data is $\{(N_{i},S_{i})\mid 1\leq i\leq n\}$, the training loss of schema router $\mathcal{M}_{r}$ with parameters $\theta_{r}$ is:
@@ 3.5. Schema Router Training & Inference
After training the schema router, we further design a graph-based constrained decoding algorithm to guarantee the validity of generated schemas. As shown in Figure 4, at each step of autoregressive decoding, we maintain a dynamic prefix tree containing the names of accessible nodes from decoded schema elements. Unlike the DFS serialization, the accessible table nodes can only be the neighbors of generated tables, not the neighbors of the database, unless no tables have been generated yet. Therefore, the available tokens (part of the node name) for each decoding step can be obtained by searching the prefix tree, prefixed with the tokens generated after the last element separator. Besides, multiple candidate schema sequences can be generated to improve the coverage of the target schema. We thus utilize diverse beam search (Vijayakumar et al., 2016), which allows for generating a more diverse set of candidate sequences.
Figure 4. An example of graph-based constrained decoding.
@@ 3.6. SQL Generation
By decoupling schema-agnostic NL2SQL into schema routing and SQL generation, DBCopilot can easily integrate with orthogonal advanced LLM-based NL2SQL solutions, no matter whether in-context prompt engineering methods (Pourreza and Rafiei, 2023; Gao et al., 2023; Guo et al., 2023; Chang and Fosler-Lussier, 2023) or specific NL2SQL LLMs (Defog.ai, 2023; Station, 2023). However, since the schema router can generate multiple schemas from different databases for an NL question, we explore three prompt strategies to select and incorporate these different database schemas for LLMs:
  • Best Schema Prompting: instructs LLMs with the highest probability schema generated by the schema router.
  • Multiple Schema Prompting: instructs LLMs using multiple candidate schemas by simply concatenating the table schemas from different sequences in the prompt.
  • Multiple Schema COT Prompting: instructs LLMs using multiple candidate schemas through chain-of-thought (COT) reasoning (Wei et al., 2022). The LLM is first asked to select the most relevant candidate schema and then generates SQL queries.
@@ 4. Experiments
In this section, we conduct experiments on schema routing and schema-agnostic NL2SQL to verify the effectiveness of DBCopilot.
@@ 4.1.1. Datasets
  • Spider (Yu et al., 2018): Spider is a widely-used cross-domain NL2SQL dataset with 10,181 questions and 5,693 unique SQL queries across 138 domains and 200 databases.
  • Bird (Li et al., 2023b): Bird is a recent challenging NL2SQL dataset designed to emphasize the importance of database values, spotlighting the challenges presented by dirty database content, external knowledge requirements, and SQL efficiency.
  • Fiben (Sen et al., 2020): Fiben is a dataset designed to evaluate the processing of complex business intelligence queries, often involving nesting and aggregation, for NL2SQL systems. It consists of 300 NL questions, corresponding to 237 different complex SQL queries on a database with 152 tables.
  • Spidersyn (Gan et al., 2021): It is a human-curated dataset designed to evaluate the robustness of NL2SQL models to synonym substitution. It builds on Spider with NL questions modified by replacing schema-related words with real-world paraphrases.
  • Spiderreal (Deng et al., 2021): This is another robustness dataset created from the dev split of the Spider dataset by manually modifying the original questions to remove the explicit mention of column names while leaving the SQL queries unchanged.
Table 2. Statistics of datasets. Asterisks (*) indicate that the original development sets are used as test sets.
Dataset Size Schema
Train Dev. Test # DB # Tables # Cols
Spider 7000 1034* 166 876 4503
Bird 9427 1534* 80 597 4337
Fiben - - 279 1 152 374
Spidersyn 7000 1034 166 876 4503
Spiderreal 508 166 876 4503
Dataset Adaptation.  These datasets are designed to evaluate NL2SQL on a single database, where each question is annotated with its target database. To approach the schema-agnostic NL2SQL, we make several adjustments. First, we remove the single-database constraint and treat all databases in the collection as potential query targets. Next, we extract metadata (tables and columns) from each SQL query using a SQL parser (Mao, 2023) and exclude any queries that cannot be parsed. Finally, we combine the target database and extracted metadata as a SQL query schema, together with the original NL question and SQL query, to form each instance ($N$, $S$, $Q$). This results in three database collections (Spider, Bird, and Fiben), with instances divided into training and test sets as in the original setup. The robustness variants of Spider share the same database collection. Table 2 summarizes the dataset statistics after adaptation.
Table 3. Schema routing performance on standard test sets. The best results are in bold and the second best are underlined.
Method Spider Bird Fiben

### v3

@@ 3.2. Schema Graph Construction
Unlike natural language expressions, a valid SQL query must follow the logical data model (Wikipedia contributors, 2023a) of the database. Therefore, it is necessary to inform the schema router about the inherent relationships and constraints between schema elements. In general, the SQL query schema $S=\langle D,T\rangle$ involved in a valid SQL query should satisfy the following conditions. First, the set of tables $T$ should be the subset of tables belonging to database $D$, since a SQL query should not reference tables that do not exist in the target database. Second, the tables in $T$ should cross-reference at least one other table with a common column (e.g., foreign key), so that the SQL query can connect them together via join or subquery clauses. Example 3.2 and 3.3 show two examples. If two tables reference different columns of the junction table, then this junction table is required in a SQL query for these two tables. However, if two tables reference the same column of another table, they can be linked implicitly.
@@
Given a natural language question “What are the names of the singers who performed in a concert in 2014?”, the corresponding SQL query is:
SELECT s.name
FROM singer_in_concert AS sc
JOIN singer AS s ON sc.singer_id = s.singer_id
JOIN concert AS c ON sc.concert_id = c.concert_id
WHERE c.year = 2014
This query involves the singer, concert, and singer_in_concert tables, where singer_in_concert is the join table with two different columns referring to singer and concert respectively.
Example 0.
Given a natural language question “Which rivers run through the state with the largest city in the US?”, the corresponding SQL query is:
SELECT river_name FROM river
WHERE traverse IN (
SELECT state_name FROM city
WHERE population = (
SELECT MAX(population) FROM city
This query involves only the city and river tables, both of which refer to the primary key of the state table.
To this end, we construct a directed graph $\mathcal{G}=\langle\mathcal{V},\mathcal{E}\rangle$ to represent the underlying schemata of all databases $\mathcal{D}$ and their tables $\mathcal{T}$. Concretely, the schema graph consists of three types of nodes $\mathcal{V}=\{\nu_{s}\}\cup\mathcal{D}\cup\mathcal{T}$ and many types of edges =$\mathcal{E}=\{e_{1},\ldots,e_{|\mathcal{E}|}\}$, where $\nu_{s}$ is a special node denoting the set of all databases $\mathcal{D}$. Each edge $e\in\mathcal{E}$ in this graph represents a relation between two nodes that can be divided into two categories:
  • Inclusion relations: Inclusion relations indicate whether a database is part of the database collection and whether a table belongs to a database.
  • Table relations: Table relations include explicit Primary-Foreign, implicit Foreign-Foreign, and Joinable relations between table nodes. Joinable (Sarma et al., 2012) indicates that two tables share some values in certain attributes, allowing them to be combined together in SQL queries.
@@
Input: A dictionary of all database schemata $databases$.
@@ 4       $links\leftarrow$ a dictionary with empty set as default value
             // Joinable covers the cases of Primary-Foreign and Foreign-Foreign.
8             for $relatedTable$ in getJoinableTables($table$) :
9                   $\mathcal{G}$.addEdge($table$, $relatedTable$))
10                   $\mathcal{G}$.addEdge($relatedTable$, $table$))
11
@@
16return $\mathcal{G}$
@@
Algorithm 1 shows the procedure for constructing this three-tiered hierarchical graph. By capturing the logical structure of database schemata in graph $\mathcal{G}$, any valid single-database SQL query schema $S$ is a trail (Wikipedia contributors, 2023c) on this graph. Each schema trail starts with node $\nu_{s}$, followed by a database node and some of its interconnected table nodes. In this way, the schema graph facilitates relation-based serialization, valid SQL query schema sampling, and graph-based constrained decoding introduced below, serving as the foundation of DBCopilot.
@@ 3.3. Schema Serialization
In this paper, we model schema routing as a generative retrieval process from an NL question $N$ to a SQL query schema $\langle D,T\rangle$. Since a Seq2Seq LM takes a sequence of tokens as input and outputs a sequence of tokens, we need to design a schema serialization algorithm that converts the partially ordered SQL query schema into a sequence of tokens.
Basic Serialization.  A trivial solution is to randomly order the set of tables as DSI for ad-hoc retrieval (Chen et al., 2023) and identify the target database based on the votes of the generated tables after inference. However, this serialization loses sight of the inclusion and table relations between schema elements. Due to the position bias and autoregressive nature of Seq2Seq LMs, the order of schema elements is important for training routers to better model their relationships (Liu et al., 2024; Berglund et al., 2024). Therefore, we need a carefully designed serialization strategy that preserves relationships within SQL query schema.
@@
where $\pi$ is the node iteration order and $\operatorname{DFS_{\pi}}(S,\mathcal{G})$ is the DFS traversal sequence of schema $S$ on schema graph $\mathcal{G}$ when iterating successor nodes under order $\pi$ and $[i]$ indicates the $i$-th node of the traversal sequence.
We introduce the iteration order parameter $\pi$ to determine the DFS sequence because a SQL query schema is not a completely ordered set and there may be multiple DFS sequences for a SQL query schema. We randomly select one DFS sequence at a time as the model target during training, and the schema router assigns different probabilities to different sequences originating from the same schema during inference.
@@
Input: An ordered pair of a SQL query schema $S=\langle D,T\rangle$, the schema graph $\mathcal{G}$, and a node iteration order $\pi$.
@@ 10      $successors$leftarrow [][\;][ ]
       // Iterates over successor nodes according to node iteration order $\pi$
@@ 3.4. Training Data Synthesis
Based on the schema graph and DFS serialization, we can train the schema router on question-schema pairs ($N$, $S$) as described in § 3.5. However, it is infeasible to manually label training data for massive database schemata. Furthermore, the generative retrieval paradigm prevents the schema router from generalizing to unseen schemata because they are not injected into model parameters. To address these challenges, we introduce a labor-free training data synthesis method that automatically generates pseudo-instances via a reverse schema-to-question generation paradigm, as illustrated in Figure 2.
Specifically, we sample a large number of valid SQL query schemata from the schema graph and generate a pseudo-question for each sampled schema. These schemata are sampled by performing finite-length random walks starting at node $\nu_{s}$ over the schema graph, and the database and tables traversed by a walk form a sampled schema. To generate an NL question for each sampled schema, we train a schema questioning model $\mathcal{M}_{\text{q}}$ in reverse based on the training sets of NL2SQL datasets. The reason for training a specific questioning model rather than using LLMs directly is to save the cost of synthesizing large amounts of training data. We use a SQL parser to extract the metadata (tables and columns) of SQL queries from the training pairs ($N$, $Q$). Unlike the schema router, which takes a condensed SQL query schema as output, the schema questioner accepts more detailed schema (e.g., table columns and comments) as input to generate rich and detailed questions. Given an illustration of the model input and output in Figure 3, the training loss of the schema questioning model $\mathcal{M}_{q}$ with parameters $\theta_{\text{q}}$ on $m$ extracted schema-question pairs ($S$, $N$) is:
@@
In practice, it can also generate high-quality seed data from schemata through LLMs and further boost them in this way.
Pseudo-Question Generation Input: Ask a question that needs to be answered by combining the contents of all the tables, based on the word1 database schema provided below. - countrylanguage(language, countrycode, official) [Optional schema comments, if available.] - country(code, lifeexpectancy) Output: What is the official language of the country with the highest life expectancy?
@@
3.5. Schema Router Training & Constrained Inference
Through the above measures, we can efficiently train the schema router to parameterize the semantic mapping of massive database schemata and effectively route NL questions to their target schemata via graph-based constrained decoding. Formally, the Seq2Seq schema router $\mathcal{M}_{r}$ takes a token sequence of question $N$ as input and generates the serialized schema sequence $\bar{S}$. Suppose the synthetic data is $\{(N_{i},S_{i})\mid 1\leq i\leq n\}$, the training loss of schema router $\mathcal{M}_{r}$ with parameters $\theta_{r}$ is:
@@ 3.5. Schema Router Training & Constrained Inference
Figure 4. An example of graph-based constrained decoding. The underline (_) denotes the separator between schema elements, and a red cross indicates that the generation of the token is invalid and disallowed. We leverage diverse beam search to ensure a more varied set of candidate sequences, represented in different colors in this graph.
After training the schema router, we further design a graph-based constrained decoding algorithm to guarantee the validity of generated schemata. As shown in Figure 4, at each step of autoregressive decoding, we maintain a dynamic prefix tree containing the names of accessible nodes from decoded schema elements. Unlike the DFS serialization, the accessible table nodes can only be the neighbors of generated tables, not the neighbors of the database, unless no tables have been generated yet. Therefore, the available tokens (i.e., part of the node name) for each decoding step can be obtained by searching the prefix tree, prefixed with the tokens generated after the last element separator. Besides, multiple candidate schema sequences can be generated to improve the coverage of the target schema. We thus utilize diverse beam search (Vijayakumar et al., 2016) to generate more diverse sets of candidate sequences and combine tables from schema sequences that share the same database to form the final candidate schemata.
@@ 3.6. SQL Generation
Basic Prompt ### Complete sqlite SQL query only and with no explanation ### Sqlite SQL tables, with their properties: # # country(code, name, continent, region, ...) # countrylanguage(countrycode, language, ...) # ### Which language is the most popular on the Asian continent? SELECT
Figure 5. An example of a basic NL2SQL prompt.
By decoupling schema-agnostic NL2SQL into schema routing and SQL generation, DBCopilot can easily integrate with orthogonal, advanced LLM-based NL2SQL solutions, no matter whether in-context prompt engineering methods (Pourreza and Rafiei, 2023; Gao et al., 2024; Guo et al., 2023; Chang and Fosler-Lussier, 2023) or specific NL2SQL LLMs (Defog.ai, 2023; Station, 2023; Li et al., 2024c). However, since the schema router can generate multiple schemata from different databases for an NL question, we explore three prompt strategies to select and incorporate these candidate schemata for LLM-based NL2SQL:
  • Best Schema Prompting: This strategy instructs LLMs with the highest probability schema generated by the schema router, as shown in Figure 5. This basic prompt template has been shown to be the most effective prompt for SQL generation with OpenAI LLMs (Gao et al., 2024).
  • Multiple Schema Prompting: This strategy instructs LLMs with multiple schemata generated by the schema router for more possible coverage of the target schema. The prompt format is the same as the best schema prompting strategy and we concatenate multiple table schemata in the prompt.
  • Multiple Schema COT Prompting: This strategy instructs LLMs using multiple candidate schemata through chain-of-thought (COT) reasoning. As shown in Figure 6, the LLM is first asked to select the most relevant candidate schema and then generates SQL queries.
Chain-of-Thought Prompt Turn 1: Based on the provided natural language question, find the database that can best answer this question from the list schemata below. Only output the corresponding database schema identifier in the [id] format, without any additional information. Question: Which language is the most popular on the Asian continent? Sqlite SQL databases, with their tables and properties: [1] car
continents(contid, continent)
countries(countryid, countryname, continent)
car_makers(id, marker, fullname, country)
[2] word
country(code, name, continent, region, ...)
countrylanguage(countrycode, language, ...)
[3] ... [4] ... [5] ...
Turn 2: Filled basic prompt with Turn 1 selected schema.
Figure 6. An example of a chain-of-thought NL2SQL prompt with multiple candidate schemata.
@@ 4. Experiments
In this section, we conduct experiments to verify the feasibility and effectiveness of DBCopilot. First, we compare DBCopilot with various retrieval-based methods for schema routing. Then, we evaluate the performance of end-to-end schema-agnostic NL2SQL. Finally, we perform detailed ablation studies.
@@ 4.1. Experimental Setup
4.1.1. Research Questions
We conducted experiments to answer the following key research questions (RQs):
RQ1: How does DBCopilot compare with various retrieval baselines in schema routing?
RQ2: Is DBCopilot more robust against semantic mismatch between NL questions and schemata?
RQ3: Does improving schema routing lead to better end-to-end schema-agnostic NL2SQL?
RQ4: What is the best prompt strategy for LLM-based SQL generation when there are multiple candidate schemata?
RQ5: How does each module (i.e., DFS serialization, data synthesis, decoding strategies) contribute to DBCopilot?
@@ 4.1.2. Datasets
  • Spider (Yu et al., 2018): Spider is a widely-used, cross-domain NL2SQL dataset with 10,181 questions and 5,693 unique SQL queries across 138 domains and 200 databases, requiring to generalize across different SQL queries and database schemas.
  • Bird (Li et al., 2023b): Bird is a challenging NL2SQL dataset designed to emphasize the importance of database content, spotlighting the challenges presented by dirty database values, external knowledge requirements, and SQL efficiency.
  • Fiben (Sen et al., 2020): Fiben is a dataset designed to simulate real-world data mart by IBM that closely mirrors actual enterprise data warehouses. Its schema conforms to standard financial ontologies, representing the complexity of schema routing in production environments. Fiben consists of 300 NL questions, corresponding to 237 complex SQL queries. These queries focus on business intelligence scenarios involving nesting and aggregation, making it an ideal testbed for evaluating NL2SQL systems in real-world scenarios.
  • Spidersyn (Gan et al., 2021): Spidersyn is a dataset designed to evaluate the robustness of NL2SQL models to synonym substitution, built on top of Spider with NL questions modified by replacing schema-related words with real-world paraphrases.
  • Spiderreal (Deng et al., 2021): Spiderreal is another robustness dataset created from the dev split of the Spider dataset by manually modifying the original NL questions to remove the explicit mention of column names.
Table 2. Statistics of datasets.
{NiceTabular}
ccccccc Dataset Size Schema
Train Dev. Test # DBs # Tables # Cols
[…]
Fiben – – 279 1 152 374
Spidersyn 7000 – 1034 166 876 4503
Spiderreal – – 508 166 876 4503
Table 3. Schema routing performance on regular test sets.
{NiceTabular}
[…]
@@
Dataset Adaptation.  These datasets are designed to evaluate NL2SQL on a single database, where each question is annotated with its target database. To approach the schema-agnostic NL2SQL, we make several adjustments. First, we remove the single-database constraint and treat all databases in the collection as potential query targets. Next, we extract metadata (tables and columns) from each SQL query using a SQL parser (Mao, 2023) and exclude any queries that cannot be parsed. Finally, we combine the target database and extracted metadata as a SQL query schema, along with the original NL question and SQL query, to form each instance ($N$, $S$, $Q$). This results in three database collections (Spider, Bird, and Fiben), with instances divided into training and test sets as in the original setup. The robustness variants of Spider share the same database collection. § 4.1.2 summarizes the dataset statistics after adaptation.